The volume of a rectangular storage box i 1440in cubed. It is 20in long, and it is half as tall as it is wide. What is the width and height of the box?
Found 2 solutions by solver91311, jim_thompson5910:
Answer by solver91311(24713)



The volume of the box is , but you know that the length is so the product of the width and the height must be divided by . But you also know that two times the height is the width, so .



Solve for the positive value of . Then calculate

John

My calculator said it, I believe it, that settles it

Answer by jim_thompson5910(35256)



"The volume of a rectangular storage box is 1440in cubed"
V = 1440

Length: L = 20
"it is half as tall as it is wide"
Width = W
Height: H = W/2

Volume of the box
V = L*W*H
V = 20*W*(W/2)
1440 = 20*W*(W/2)
1440 = 10*W^2
10*W^2 = 1440
W^2 = 1440/10
W^2 = 144
W = sqrt(144)
W = 12

H = W/2
H = 12/2
H = 6

Width = 12 inches
Height = 6 inches